OpenStudy (anonymous):

Determine the better investment: 5% compounded monthly or 5.2% compounded semiannually

OpenStudy (anonymous):

It's like this, no?\[ \left(1+\frac{0.05}{12}\right)^{12} \]\[ \left(1+\frac{0.05}{2}\right)^{2} \]

OpenStudy (anonymous):

so, which is better?

OpenStudy (anonymous):

Put it in calculator.

OpenStudy (anonymous):

I should have put: \[ \left(1+\frac{0.05}{12}\right)^{12} \\ \left(1+\frac{0.052}{2}\right)^{2} \]